# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ Appendix D Stochastic Implicit Differentiation
Thanks to the chain rule of conservative derivatives, if 5.1(i) holds and we set, for every $u\in\mathbb{R}^{d},\lambda\in O_{\Lambda}$
$D_{\Phi}(u,\lambda)=D_{\Phi,1}(u,\lambda)\times D_{\Phi,2}(u,\lambda)$
with
$D_{\Phi,1}(u,\lambda)$ $=D_{G,1}(T(u,\lambda),\lambda)D_{T,1}(u,\lambda)$
$D_{\Phi,2}(u,\lambda)$ $=D_{G,1}(T(u,\lambda),\lambda)D_{T,2}(u,\lambda)+D_{G,2}(T(u,\lambda),\lambda),$
then $D_{\Phi}$ is a conservative derivative for $\Phi$. Furthermore if 5.1(i) is satisfied, then $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$ and $D_{w}^{\operatorname{fix}}$ and $D_{w}^{\operatorname{imp}}$ in (11) and (10) are well defined and conservative derivatives of $w$.
The following result is similar to Lemma 3.2 and follows again from Lemma 2.5. The only difference is that the constants are majorized so to be independent on $u$. This is done only to simplify the analysis.
@@
Under 5.1, for every $\lambda\in\Lambda$, there exist $R_{G,\lambda},R_{T,\lambda}>0$ such that for every $u\in\mathbb{R}^{d}$ and $k\in\{1,2\}$
$\operatorname{\it{e}}(D_{G,k}(u,\lambda),D_{G,k}(T(w(\lambda),\lambda),\lambda)$ $\leq C_{G,\lambda}\lVert u-T(w(\lambda),\lambda)\rVert$
$\operatorname{\it{e}}(D_{T,k}(u,\lambda),D_{T,k}(w(\lambda),\lambda))$ $\leq C_{T,\lambda}\lVert u-w(\lambda)\rVert,$
where
$C_{G,\lambda}$ $:=L_{G}+M_{G,\lambda}/R_{G,\lambda}\quad C_{T,\lambda}:=L_{T}+M_{T,\lambda}/R_{T,\lambda}$ (24)
with $M_{X,\lambda}:=\max_{i\in\{1,\dots,m\}}\min_{j\in I_{X}(w(\lambda),\lambda)}\lVert X^{\prime}_{i}(w(\lambda),\lambda)-X^{\prime}_{j}(w(\lambda),\lambda)\rVert$ and $L_{X},I_{X}$ satisfy 3.1(i) and $X\in\{G,T\}$.
Proof.
We proceed similarly to the proof of Lemma 3.2 with $\Phi$ replaced by $T$ and $G$ and, in the case of $G$ we also replace $w(\lambda)$ with $T(w(\lambda),\lambda)$. ∎
@@
Set, for the sake of brevity, $z_{t}=(w_{t}(\lambda),\lambda)$ and $z=(w(\lambda),\lambda)$. We also set $v_{k}=v_{k}(w_{t}(\lambda),\bar{T}_{t}(\lambda))$ and $\bar{v}=\bar{v}(w_{t}(\lambda),\bar{T}_{t}(\lambda))$.
Let $\Delta_{\lambda}=\lVert w(\lambda)-T(z)\rVert$. Then from Assumption 5.2 on the variance of $\hat{T}$ and since $T(\cdot,\lambda)$ is $1$-Lipschitz we have
$\operatorname{Var}[\hat{T}(z_{t},\xi)\,|\,w_{t}(\lambda),y]$ $\leq\sigma_{1}+\sigma_{2}\lVert w_{t}(\lambda)\mp w(\lambda)-T(w_{t}(\lambda),\lambda)\pm T(w(\lambda),\lambda)\rVert^{2}$
$\leq\sigma_{1}+3\sigma_{2}(2\Delta_{t}^{2}+\Delta_{\lambda}^{2})=:B_{3}(\Delta_{t}^{2}).$ (25)
Using the properties of the excess in Lemma B.1 and recalling that
$D_{w}^{\operatorname{imp}}(\lambda)=(I-D_{\Phi,1}(z))^{-1}D_{\Phi,2}(z),$
we have the following error decomposition
$\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y)$ $\leq\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k},D_{\Phi,2}(z)^{\top}v_{k})+\operatorname{\it{e}}(D_{\Phi,2}(z)^{\top}v_{k},D_{\Phi,2}(z)^{\top}(I-D_{\Phi,1}(z)^{\top})^{-1}y)$
$\leq\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t}),D_{\Phi,2}(z))\lVert v_{k}\rVert+\lVert D_{\Phi,2}(z)\rVert_{\sup}\operatorname{\it{e}}(v_{k},(I-D_{\Phi,1}(z)^{\top})^{-1}y),$
and hence, squaring and taking the expectation of both sides and using the law of total expectation yields
$\mathbb{E}[\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y)^{2}]\,|\,w_{t}(\lambda),y]$ $\leq\underbrace{2\mathbb{E}\big{[}\mathbb{E}[\lVert v_{k}\rVert^{2}\,|\,w_{t}(\lambda),y,\bm{\xi}^{(1)}]\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t}),D_{\Phi,2}(z))^{2}\,\big{|}\,w_{t}(\lambda),y\big{]}}_{(1)}$ (26)
$\qquad+\underbrace{2\lVert D_{\Phi,2}(z)\rVert^{2}_{\sup}\mathbb{E}[\operatorname{\it{e}}(v_{k},(I-D_{\Phi,1}(z)^{\top})^{-1}y)^{2}\,|\,w_{t}(\lambda),y]}_{(2)}.$
Bound for the first term in the RHS of (26)
@@
where in the second last inequality we used 5.3(ii). Note that the bound above is independent from $\bm{\xi}^{(1)}$ and hence
$(1)\leq 4\lVert y\rVert^{2}\big{(}\sigma_{\lambda}(k)+1/(1-q)^{2}\big{)}\mathbb{E}[\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t}),D_{\Phi,2}(z))^{2}\,|\,w_{t}(\lambda),y].$
@@
$\partial_{2}\bar{\Phi}(z_{t})$ $=\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{2}\bar{T}(z_{t})+\partial_{2}G(\bar{T}(z_{t}),\lambda)$
$D_{\Phi,2}(z)$ $=D_{G,1}(T(z),\lambda)D_{T,2}(z)+D_{G,2}(T(z),\lambda),$
therefore using the properties of excess in Lemma B.1 we have
$\operatorname{\it{e}}\!\big{(}$ $\partial_{2}\bar{\Phi}(z_{t}),D_{\Phi,2}(z)\big{)}$
$\leq\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{2}\bar{T}(z_{t}),\partial_{1}G(\bar{T}(z_{t}),\lambda)D_{T,2}(z)\big{)}$
$\qquad+\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda)D_{T,2}(z),D_{G,1}(T(z),\lambda)D_{T,2}(z)\big{)}+\operatorname{\it{e}}\!\big{(}\partial_{2}G(\bar{T}(z_{t}),\lambda),D_{G,2}(T(z),\lambda)\big{)}$
$\leq\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\rVert\operatorname{\it{e}}(\partial_{2}\bar{T}(z_{t}),D_{T,2}(z))$
$\qquad+\lVert D_{T,2}(z)\rVert_{\sup}\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda),D_{G,1}(T(z),\lambda)\big{)}+\operatorname{\it{e}}\!\big{(}\partial_{2}G(\bar{T}(z_{t}),\lambda),D_{G,2}(T(z),\lambda)\big{)}$
$\leq\lVert\partial_{2}\bar{T}(z_{t})-\partial_{2}T(z_{t})\rVert+\operatorname{\it{e}}(\partial_{2}T(z_{t}),D_{T,2}(z))$
$\qquad+\lVert D_{T,2}(z)\rVert_{\sup}\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda),D_{G,1}(T(z),\lambda)\big{)}+\operatorname{\it{e}}\!\big{(}\partial_{2}G(\bar{T}(z_{t}),\lambda),D_{G,2}(T(z),\lambda)\big{)}$
$\leq\lVert\partial_{2}\bar{T}(z_{t})-\partial_{2}T(z_{t})\rVert+C_{T,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert$
$\qquad+C_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})(\lVert\bar{T}(z_{t})-T(z_{t})\rVert+\lVert T(z_{t})-T(z)\rVert)$
@@
$\qquad+C_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})\lVert\bar{T}(z_{t})-T(z_{t})\rVert.$
Hence, recalling Assumption 5.2 and (D), we have
$\mathbb{E}\big{[}\!\operatorname{\it{e}}\!\big{(}\partial_{2}\bar{\Phi}(z_{t}),D_{\Phi,2}(z)\big{)}^{2}\,\big{|}\,w_{t}(\lambda),y\big{]}$ $\leq 3\operatorname{Var}[\partial_{2}\bar{T}(z_{t})\,|\,w_{t}(\lambda),y]$
$\qquad+3[C_{T,\lambda}+C_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})]^{2}\Delta_{t}^{2}$
@@
$\qquad+3C^{2}_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})^{2}\frac{B_{3}(\Delta_{t}^{2})}{J}.$
@@
$(1)\leq 12\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{2}}\bigg{)}\bigg{(}\frac{\sigma_{2}^{\prime}}{J}+[C_{T,\lambda}+C_{G,\lambda}M_{T,\lambda}]^{2}\Delta_{t}^{2}+C^{2}_{G,\lambda}M_{T,\lambda}^{2}\frac{B_{3}(\Delta_{t}^{2})}{J}\bigg{)},$
@@
$\operatorname{\it{e}}(v_{k},(I-D_{\Phi,1}(z)^{\top})^{-1}y)\leq\lVert v_{k}-\bar{v}\rVert+\operatorname{\it{e}}(\bar{v},(I-D_{\Phi,1}(z))^{-\top}y).$
Now we recall that $\bar{v}=\big{(}I-\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t})\big{)}^{-\top}$. Hence, setting $A=I-\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t})$ and $\mathcal{B}=I-D_{\Phi,1}(z)$, and using the properties of Lemma B.1(iv)(v), Lemma D.1, and noting that $\lVert\partial_{1}T(z_{t})\rVert,\lVert D_{G,1}(z)\rVert_{\sup}\leq 1$ and $\lVert A^{-1}\rVert,\lVert\mathcal{B}^{-1}\rVert_{\sup}\leq 1/(1-q)$ (Lemma B.3) we obtain
$\operatorname{\it{e}}(v_{k},(I$ $-D_{\Phi,1}(z)^{\top})^{-1}y)$
$\leq\lVert v_{k}-\bar{v}\rVert+\lVert y\rVert\lVert A^{-1}\rVert\lVert\mathcal{B}^{-1}\rVert_{\sup}\operatorname{\it{e}}(A,\mathcal{B})$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t}),D_{G,1}(T(z),\lambda)D_{T,1}(z)\big{)}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t}),D_{G,1}(T(z),\lambda)\partial_{1}T(z_{t})\big{)}$
$\qquad\qquad+\operatorname{\it{e}}\!\big{(}D_{G,1}(T(z),\lambda)\partial_{1}T(z_{t}),D_{G,1}(T(z),\lambda)D_{T,1}(z)\big{)}\Big{]}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\lVert\partial_{1}T(z_{t})\rVert\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda),D_{G,1}(T(z),\lambda)\big{)}$
$\qquad\qquad+\lVert D_{G,1}(T(z),\lambda)\rVert_{\sup}\operatorname{\it{e}}\!\big{(}\partial_{1}T(z_{t}),D_{T,1}(z)\big{)}\Big{]}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\operatorname{\it{e}}\!\big{(}\partial_{1}G(\bar{T}(z_{t}),\lambda),D_{G,1}(T(z),\lambda)\big{)}+\operatorname{\it{e}}\!\big{(}\partial_{1}T(z_{t}),D_{T,1}(z)\big{)}\Big{]}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\big{[}C_{G,\lambda}(\lVert\bar{T}(z_{t})-T(z_{t})\rVert+\lVert T(z_{t})-T(z)\rVert)+C_{T,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert\big{]}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\big{[}C_{G,\lambda}\lVert\bar{T}(z_{t})-T(z_{t})\rVert+(C_{G,\lambda}+C_{T,\lambda})\Delta_{t}\big{]}.$
Therefore,
$\mathbb{E}\big{[}\!\operatorname{\it{e}}(v_{k},$ $(I-D_{\Phi,1}(z)^{\top})^{-1}y)^{2}\,\big{|}\,w_{t}(\lambda),y,\bm{\xi}^{(1)}\big{]}$
@@
and hence
$\mathbb{E}\big{[}\!\operatorname{\it{e}}(v_{k},$ $(I-D_{\Phi,1}(z)^{\top})^{-1}y)^{2}\,\big{|}\,w_{t}(\lambda),y\big{]}$

### v2

@@ Appendix D Stochastic Implicit Differentiation
For simplicity let for every $u\in\mathbb{R}^{d},\lambda\in O_{\Lambda}$
$\Psi(u,\lambda)=(T(u,\lambda),\lambda),\qquad D_{\Psi}(u,\lambda)=\left\{\begin{bmatrix}C_{1}&C_{2}\\ 0&I_{m}\end{bmatrix}\,\Big{|}\,[C_{1},C_{2}]\in D_{T}(u,\lambda)\right\}$
From Lemma 3 in (Bolte & Pauwels, 2021) and 5.1(i) it follows that $D_{\Psi}$ is a conservative derivative of $\Psi$. Moreover, we can write $\Phi(u,\lambda)=G(\Psi(u,\lambda))$ and thanks to the chain rule of conservative derivatives we have that
$D_{\Phi}(u,\lambda):=D_{G}(\Psi(u,\lambda))D_{\Psi}(u,\lambda)=D_{G}(T(u,\lambda),\lambda)\begin{bmatrix}D_{T}(u,\lambda)\\ 0\ \ I_{m}\end{bmatrix}$ (24)
is a conservative derivative for $\Phi$. Furthermore, if 5.1(ii) is satisfied, then $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$ and $D_{w}^{\operatorname{fix}}$ and $D_{w}^{\operatorname{imp}}$ in (12) and (11) are well defined and conservative derivatives of $w$. Similarly, a conservative derivative of $\bar{\Phi}$ can be obtained as
$D_{\bar{\Phi}}(u,\lambda):=D_{G}(\bar{T}(u,\lambda),\lambda)\begin{bmatrix}D_{\bar{T}}(u,\lambda)\\ 0\ \ I_{m}\end{bmatrix},\quad with\quad D_{\bar{T}}(u,\lambda)=\frac{1}{J}\sum_{j=1}^{J}D_{\hat{T}_{\xi^{(1)}_{j}}}(u,\lambda).$ (25)
Note that $\partial_{2}\bar{\Phi}(u,\lambda)$ as defined in (21) is an element of $D_{\bar{\Phi},2}$.
The following result is similar to Lemma 3.2 and follows directly from Lemma 2.5. The only difference is that the constants are majorized so to be independent on $u$. This is done only to simplify the analysis.
@@
Under 5.1, for every $\lambda\in\Lambda$, there exist $R_{G,\lambda},R_{T,\lambda}>0$ such that for every $u\in\mathbb{R}^{d}$ and
$\operatorname{\it{e}}(D_{G}(u,\lambda),D_{G}(T(w(\lambda),\lambda),\lambda)$ $\leq C_{G,\lambda}\lVert u-T(w(\lambda),\lambda)\rVert$
$\operatorname{\it{e}}(D_{T}(u,\lambda),D_{T}(w(\lambda),\lambda))$ $\leq C_{T,\lambda}\lVert u-w(\lambda)\rVert,$
where $C_{G,\lambda}:=L_{G}+M_{G,\lambda}/R_{G,\lambda}$, $C_{T,\lambda}:=L_{T}+M_{T,\lambda}/R_{T,\lambda}$, with
$M_{T,\lambda}$ $:=\max_{i\in\{1,\dots,r\}}\min_{j\in I_{T}(w(\lambda),\lambda)}\lVert T^{\prime}_{i}(w(\lambda),\lambda)-T^{\prime}_{j}(w(\lambda),\lambda)\rVert$
$M_{G,\lambda}$ $:=\max_{i\in\{1,\dots,r\}}\min_{j\in I_{G}(T(w(\lambda),\lambda),\lambda)}\lVert G^{\prime}_{i}(T(w(\lambda),\lambda),\lambda)-G^{\prime}_{j}(T(w(\lambda),\lambda),\lambda)\rVert$
and $L_{T},L_{G}$ satisfying 3.1(i).
@@
Set, for the sake of brevity, $z_{t}=(w_{t}(\lambda),\lambda)$ and $z=(w(\lambda),\lambda)$. We also set $v_{k}=v_{k}(w_{t}(\lambda),\bar{T}_{t}(\lambda))$, $\bar{v}=\bar{v}(w_{t}(\lambda),\bar{T}_{t}(\lambda))$, and $a_{\lambda}=\lVert w(\lambda)-T(z)\rVert$. From Assumption 5.2 on the variance of $\hat{T}$ and since $T(\cdot,\lambda)$ is $1$-Lipschitz we have
$\operatorname{Var}[\hat{T}_{\xi}(z_{t})\,|\,w_{t}(\lambda),y]$ $\leq\sigma_{1}+\sigma_{2}\lVert w_{t}(\lambda)\mp w(\lambda)-T(z_{t})\pm T(z)\rVert^{2}$
$\leq\sigma_{1}+3\sigma_{2}(2\Delta_{t}^{2}+a_{\lambda}^{2})=:b_{\lambda}(\Delta_{t}^{2}).$ (26)
Now, recall that $G^{\prime}(\bar{T}(z_{t}),\lambda)=[\partial_{1}G(\bar{T}(z_{t}),\lambda),\partial_{2}G(\bar{T}(z_{t}),\lambda)]$, $T^{\prime}(z_{t})=[\partial_{1}T(z_{t}),\partial_{2}T(z_{t})]$ and set
$B$ $:=[B_{1},B_{2}]=\operatorname*{arg\,min}_{B^{\prime}\in D_{G}(T(z),\lambda)}\lVert G^{\prime}(\bar{T}(z_{t}),\lambda)-B^{\prime}\rVert$
$C$ $:=[C_{1},C_{2}]=\operatorname*{arg\,min}_{C^{\prime}\in D_{T}(z)}\lVert T^{\prime}(z_{t})-C^{\prime}\rVert$
with $B_{1},C_{1}\in\mathbb{R}^{d\times d}$, $B_{2},C_{2}\in\mathbb{R}^{d\times m}$, which is valid since the $\operatorname*{arg\,min}$ is over compact convex sets. Then, recalling the definition of excess and applying Lemma D.1 we have that for $j=1,2$
$\lVert\partial_{j}G(\bar{T}(z_{t}),\lambda)-B_{j}\rVert$ $\leq\lVert G^{\prime}(\bar{T}(z_{t}),\lambda)-B\rVert=\operatorname{\it{e}}(G^{\prime}(\bar{T}(z_{t}),\lambda),D_{G}(T(z),\lambda))\leq C_{G,\lambda}\lVert\bar{T}(z_{t})-T(z)\rVert,$ (27)
$\lVert\partial_{j}T(z_{t})-C_{j}\rVert$ $\leq\lVert T^{\prime}(z_{t})-C\rVert=\operatorname{\it{e}}(T^{\prime}(z_{t}),D_{G}(z))\leq C_{G,\lambda}\lVert z_{t}-z\rVert.$
Let also $A:=[A_{1},A_{2}]$ with $A_{1}:=B_{1}C_{1}$ and $A_{2}:=B_{1}C_{2}+B_{2}$. Since $B\in D_{G}(T(z),\lambda)$, $C\in D_{T}(z)$ we have that $A\in D_{\Phi}(z)$ (from the definition of $D_{\Phi}$ in (24)) and consequently that $(I-A_{1})^{-1}A_{2}\in D_{w}^{\operatorname{imp}}(\lambda)$. Hence, recalling the definition of excess we can write
$\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y)\leq\lVert\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k}-A_{2}^{\top}(I-A_{1})^{-\top}y\rVert.$
To prove the result, it is therefore sufficient to appropriately control the distance to a particular element of $D_{w}^{\operatorname{imp}}(\lambda)^{\top}y$, namely $A_{2}^{\top}(I-A_{1})^{-\top}y$, which is a random variable depending on $y,w_{t}(\lambda),\bm{\xi}^{(1)}$ (from the definition of $B$ and $C$). We have the following error decomposition
$\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k},D_{w}^{\operatorname{imp}}(\lambda)^{\top}y)$ $\leq\lVert\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k}-A_{2}^{\top}v_{k}\rVert+\lVert A_{2}^{\top}v_{k}-A_{2}^{\top}(I-A_{1}^{\top})^{-1}y\rVert)$
$\leq\lVert\partial_{2}\bar{\Phi}(z_{t})-A_{2}\rVert\lVert v_{k}\rVert+\lVert D_{\Phi,2}(z)\rVert_{\sup}\lVert v_{k}-(I-A_{1}^{\top})^{-1}y\rVert,$
where we used that $A_{2}\in D_{\Phi,2}(z)$. Hence, squaring and taking the conditional expectation of both sides yields
$\mathbb{E}[\operatorname{\it{e}}(\partial_{2}\bar{\Phi}(z_{t})^{\top}v_{k},D_{w}^{\operatorname{imp}}(\lambda)^{\top}$ $y)^{2}]\,|\,w_{t}(\lambda),y,\bm{\xi}^{(1)}]\leq\underbrace{2\mathbb{E}[\lVert v_{k}\rVert^{2}\,|\,w_{t}(\lambda),y,\bm{\xi}^{(1)}]\cdot\lVert\partial_{2}\bar{\Phi}(z_{t})-A_{2}\rVert^{2}}_{(1)}$ (28)
$\quad+\underbrace{2\lVert D_{\Phi,2}(z)\rVert_{\sup}^{2}\mathbb{E}[\lVert v_{k}-(I-A_{1}^{\top})^{-1}y\rVert^{2}\,|\,w_{t}(\lambda),y,\bm{\xi}^{(1)}]}_{(2)}.$
Bound for term (1) of (28)
@@
where in the second last inequality we used 5.4(ii). Hence
$(1)\leq 4\lVert y\rVert^{2}\big{(}\sigma_{\lambda}(k)+1/(1-q)^{2}\big{)}\lVert\partial_{2}\bar{\Phi}(z_{t})-A_{2}\rVert^{2}.$
@@
$\partial_{2}\bar{\Phi}(z_{t})$ $=\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{2}\bar{T}(z_{t})+\partial_{2}G(\bar{T}(z_{t}),\lambda),\quad A_{2}=B_{1}C_{2}+B_{2},$
therefore we have
$\lVert\partial_{2}\bar{\Phi}(z_{t})-A_{2}\rVert$ $\leq\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{2}\bar{T}(z_{t})-\partial_{1}G(\bar{T}(z_{t}),\lambda)C_{2}\rVert$
$\qquad+\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)C_{2}-B_{1}C_{2}\rVert+\lVert\partial_{2}G(\bar{T}(z_{t}),\lambda)-B_{2}\rVert$
$\leq\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\rVert\lVert\partial_{2}\bar{T}(z_{t})-C_{2}\rVert$
$\qquad+\lVert C_{2}\rVert\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)-B_{1}\rVert+\lVert\partial_{2}G(\bar{T}(z_{t}),\lambda)-B_{2}\rVert$
$\leq\lVert\partial_{2}\bar{T}(z_{t})-\partial_{2}T(z_{t})\rVert+\lVert\partial_{2}T(z_{t})-C_{2}\rVert$
$\qquad+\lVert B_{2}\rVert\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)-B_{1}\rVert+\lVert\partial_{2}G(\bar{T}(z_{t}),\lambda)-B_{2}\rVert$
$\overset{(*)}{\leq}\lVert\partial_{2}\bar{T}(z_{t})-\partial_{2}T(z_{t})\rVert+C_{T,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert$
$\qquad+C_{G,\lambda}(1+\lVert C_{2}\rVert)(\lVert\bar{T}(z_{t})-T(z_{t})\rVert+\lVert T(z_{t})-T(z)\rVert)$
@@
$\qquad+C_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})\lVert\bar{T}(z_{t})-T(z_{t})\rVert,$
[…]
$\mathbb{E}\big{[}\lVert\partial_{2}\bar{\Phi}(z_{t})-C_{2}\big{\rVert}^{2}\,\big{|}\,w_{t}(\lambda),y\big{]}$ $\leq 3\operatorname{Var}[\partial_{2}\bar{T}(z_{t})\,|\,w_{t}(\lambda),y]+3[C_{T,\lambda}+C_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})]^{2}\Delta_{t}^{2}$
@@
$\qquad+3C^{2}_{G,\lambda}(1+\lVert D_{T,2}(z)\rVert_{\sup})^{2}\frac{b_{\lambda}(\Delta_{t}^{2})}{J}.$
@@
$\mathbb{E}[(1)\,|\,w_{t}(\lambda),y]\leq 12\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{2}}\bigg{)}\bigg{(}\frac{\sigma_{2}^{\prime}}{J}+[C_{T,\lambda}+C_{G,\lambda}M_{T,\lambda}]^{2}\Delta_{t}^{2}+C^{2}_{G,\lambda}M_{T,\lambda}^{2}\frac{b_{\lambda}(\Delta_{t}^{2})}{J}\bigg{)},$
@@
$\lVert v_{k}-(I-A_{1}^{\top})^{-1}y\rVert\leq\lVert v_{k}-\bar{v}\rVert+\lVert\bar{v}-(I-A_{1})^{-\top}y\rVert.$
Let $\hat{B}_{1}=\partial_{1}G(\bar{T}(z_{t}),\lambda)$ and $\hat{C}_{1}=\partial_{1}T(z_{t})$ and $\hat{A}_{1}=\hat{B}_{1}\hat{C}_{1}$ and recall that $\bar{v}=(I-\hat{A}_{1}^{\top})^{-1}y$ and $A_{1}\in D_{\Phi,1}(z)$. Noting that $\max\{\lVert B_{1}\rVert,\lVert C_{1}\rVert,\lVert\hat{B}_{1}\rVert,\lVert\hat{C}_{1}\rVert)\}\leq 1$, $\max\{\lVert\hat{A}_{1}\rVert,\lVert A_{1}\rVert\}\leq q$ and hence $\max\{\lVert(I-\hat{A}_{1}^{\top})^{-1}\rVert,\lVert(I-A_{1}^{\top})^{-1}\rVert\}\leq 1/(1-q)$, we obtain
$\lVert v_{k}$ $-(I-A_{1}^{\top})^{-1}y\rVert$
$\leq\lVert v_{k}-\bar{v}\rVert+\lVert y\rVert\lVert(I-\hat{A}_{1}^{\top})^{-1}\rVert\lVert(I-A_{1}^{\top})^{-1}\rVert\lVert\hat{A}_{1}-A_{1}\rVert$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t})-B_{1}C_{1}\rVert$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t})-B_{1}\partial_{1}T(z_{t})\rVert+\lVert B_{1}\partial_{1}T(z_{t})-B_{1}C_{1}\big{)}\rVert$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\lVert\partial_{1}T(z_{t})\rVert\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)-B_{1}\rVert+\lVert B_{1}\rVert\lVert\partial_{1}T(z_{t})-C_{1}\rVert\Big{]}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)-B_{1}\rVert+\lVert\partial_{1}T(z_{t})-C_{1}\rVert\Big{]}$
$\overset{(*)}{\leq}\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\big{[}C_{G,\lambda}(\lVert\bar{T}(z_{t})-T(z_{t})\rVert+\lVert T(z_{t})-T(z)\rVert)+C_{T,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert\big{]}$
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\big{[}C_{G,\lambda}\lVert\bar{T}(z_{t})-T(z_{t})\rVert+(C_{G,\lambda}+C_{T,\lambda})\Delta_{t}\big{]},$
[…]
$\mathbb{E}\big{[}\!\operatorname{\it{e}}(v_{k},$ $(I-A_{1}^{\top})^{-1}y)^{2}\,\big{|}\,w_{t}(\lambda),y,\bm{\xi}^{(1)}\big{]}$
@@
and hence, taking the expectation over $\bm{\xi}^{(1)}$ we obtain
$\mathbb{E}\big{[}\!\operatorname{\it{e}}(v_{k},$ $(I-A_{1}^{\top})^{-1}y)^{2}\,\big{|}\,w_{t}(\lambda),y\big{]}$